Solve the initial value problem y^{2}y -y^{3}tanx=sinxcos^{2}x,y(0)=1


1
Expert's answer
2021-08-26T01:05:35-0400

Given IVP:

"y^2y'-y^3\tan x=\sin x\cos^2x,\quad y(0)=1"

Let us denote

"z(x)=y^3(x)"

We get new IVP

"\frac{1}{3}z'-z\tan x=\sin x\cos^2x,\quad z(0)=1"

Let

"z=u\cdot v"

Then

"\frac{1}{3}(u'v+uv')-uv\tan x=\sin x\cos^2x"

Let

"\frac{1}{3}v'-v\tan x=0"

We obtain

"\frac{dv}{v}=3\tan x dx"

"\ln v=-3\ln \cos x=\ln \cos^{-3}x"

"v=\cos^{-3}x""\frac{1}{3}u'v=\sin x\cos^2x"

"\frac{1}{3}u'\cos^{-3}x=\sin x\cos^2x"

"u'=3\sin x\cos^5x"

"u=-\frac{\cos^6x}{2}+C"

"z(x)=u\cdot v=(-\frac{\cos^6x}{2}+C)\cos^{-3}x"

The initial condition gives

"z(0)=(-1/2+C)=1,\quad C=3/2"

So

"z(x)=\left(-\frac{\cos^6x}{2}+\frac{3}{2}\right)\cos^{-3}x"

Finally

"y(x)=\sqrt[3]{\left(-\frac{\cos^6x}{2}+\frac{3}{2}\right)\cos^{-3}x}"

"y(x)=\sqrt[3]{\left(-\frac{\cos^3x}{2}+\frac{3}{2\cos^3x}\right)}"